# SimCity: Multi-Agent Urban Development Simulation with Rich Interactions

### v1

@@ 1 Introduction
The rapid development of Large Language Models (LLMs), has enabled multi-agent simulations of human societal activities across diverse scales and domains (Gao et al., 2024a). In these settings, autonomous LLM-powered agents interact with each other and with their environment. Existing work has examined not only general social simulacra (Park et al., 2023; Huang et al., 2025; Piao et al., 2025), but also domain-specific applications such as public administration crisis (Xiao et al., 2023), health policy (Hou et al., 2025) or deduction game (Xu et al., 2025). While these work demonstrate the social simulation capabilities of LLM-driven agents, we focus on their application to the simulation and evaluation of urban-style economic activities.
For the past two decades, the Dynamic Stochastic General Equilibrium (DSGE) framework has been the predominant paradigm for studying aggregate economic behavior (Blanchard, 2009; Glandon et al., 2023). While mathematically elegant, DSGE models the economy with fully rational agents solving explicit optimization problems in equilibrium, which restricts heterogeneity and richer behavioral dynamics (Sergi, 2018; Vines & Wills, 2020; Storm, 2021). agent-based models (ABMs) offer a bottom-up alternative that accommodates heterogeneity but typically rely on hand-crafted decision rules (Dilaver et al., 2018). Advancing the ABM tradition with recent developments in LLMs, we propose SimCity, a multi-agent macroeconomic simulation framework that enables flexible, adaptive behavior with transparent natural language reasoning, and a virtual city environment that provides spatial context.
SimCity models the economy as the interaction of four types of agents: households, firms, a central bank, and a government. To incorporate realistic heterogeneity, households and firms are instantiated as families of agents with rich variation in preferences, abilities and other background, whereas the central bank and government are modeled as single institutional agents, in line with the economic literature (Blanchard, 2025). Agents are implemented through an LLM-based module that integrates environmental observations, agent-specific traits, and structured memory to generate reasoning, planning, and decision-making. Following the classic agent–interaction–environment paradigm of multi-agent system design (Wooldridge, 2009), we construct a simulation environment that features a frictional labor market, heterogeneous goods markets, and core financial interactions. To provide spatial context for both agents and researchers, the environment is augmented with a visualized map.
@@
Experimental results show that SimCity successfully exhibits a range of classic macroeconomic phenomena, such as Okun’s law, the Phillips curve, the Beveridge curve, price elasticity of demand and Engel’s curve. It also simulate the dynamic expansion of a city, and facilitates the simulation of novel economic shocks that were difficult to study in previous frameworks.
@@
1. LLM-driven macroeconomic simulation with rich interactions. We integrate LLMs into economic simulation by modeling four distinct economic roles, households, firms, a central bank, and a government, as LLM-driven agents, enabling realistic agent heterogeneity and rich economic interactions.
@@
Table 1: A comparison of agent-driven macroeconomic simulators.
Simulator AI Economist LEN CATS EconAgent Simcity
Households 10 100 100 200 200
@@
Value-Added
Agents’ Types [^1] H: Household, F: Firm, B: Bank, G: Government HG HB HFB HBG HFBG
Markets [^2] L: Labor Market, G: Goods Market, F: Financial Market $\times$ L LGF $\times$ LGF
@@
2 Related Work
LLMs for Social Simulation
Traditional Macroeconomic Modeling
DSGE and ABM represent two leading, and in many respects, contrasting approaches to macroeconomic modeling. DSGE features highly idealized decision-makers characterized by precise mathematical objective functions, rational expectations about future variables, and strict market-clearing equilibrium conditions (Stokey et al., 1989; Ljungqvist & Sargent, 2018). In contrast, ABM follows a bottom-up, simulation-driven methodology, where agents are endowed with rule-of-thumb behavior patterns and, in some cases, learning heuristics (Tesfatsion & Judd, 2006; Axtell & Farmer, 2025). Although DSGE remains the dominant approach in economics departments and policy institutions (Woodford, 2009; Negro et al., 2013; Chen et al., 2023), there has been a renewal of interests in ABMs, particularly following the Global Financial Crisis of 2008 and again after the COVID-19 crisis, as mainstream DSGE models struggled to account for these unprecedented shocks (Gatti et al., 2008; Stiglitz, 2018; Borsos et al., 2024). Our work is closer to the ABM tradition. However, we depart from conventional ABMs by replacing predetermined behavioral rules with LLM-powered agents capable of flexible reasoning and information processing in natural languages.
Closely Related Work
Li et al. (2024) introduces EconAgent as an recent attempt at LLM-driven macroeconomic simulation. Its framework restricts agent-environment interactions to two simplified variables: consumption share and work propensity, which constrains the range of economic activities that can be simulated. Our work extends this approach by incorporating other key economic elements, including firms as LLM-driven agents, heterogeneous goods markets, enriched financial activities and taxation system. Thus, our framework allows the exploration of a wider range of macroeconomic phenomena. Another contemporary work Mi et al. (2025), which build similar cluster of agents, is proposed as a testbed for performance evaluation while we focus on urban development and macroeconomic laws.
@@ 3 Environment and Interaction
Our framework consists of three core layers: environment, interaction protocol, and agents. This section describes environment and interaction of our system architecture.
@@ 3.1 Environment
Each simulation step represents one month. The order of events within each step is detailed in Appendix D.2. By design, agents can only access information from previous steps.
The simulation proceeds in two phases. In phase 1 (the move-in phase), new households with synthetic profiles, representing immigrants or newborns, are introduced into the SimCity environment until a predetermined maximum population is reached. Unlike the typical burn-in period used purely for model stabilization in ABMs (Lengnick, 2013), this phase is explicitly designed to study urban expansion, capturing how economic activity and spatial development evolve as the city grows.
In phase 2 (the development phase), the population is fixed. However urban developments such as firm creation, investment, and spatial reallocation, continue to take place. We focus on this phase to study macroeconomic regularities, as the system operates in a quasi-steady state that is less susceptible to the transient noise present during the expansion period.
@@
We develop a web-based render module to visualize the urban expansion as shown in Figure 1 shown. All buildings are displayed on the rendered map. Geographic placement of new firms are decided by a Visual-Language Model (VLM). For technical details of the module please refer to Appendix D.3.
@@ 3.2 Interaction Protocol
We leverage the common-sense reasoning capabilities of LLMs to act as human-like, heterogeneous agents. Details about the prompts used are provided in Appendix E.1. Agents interacts with the environment by means of function calling. The framework loads all the operations that the agent can execute and appends formatted function names along with their descriptions to the prompt. The LLMs will return the actions to be taken and their parameters in JSON format. Framework will execute after a verification. For detailed examples, see the Appendix E.3.
@@ 4 Agents
This section presents an overview of the agents in SimCity. As illustrated in Figure 1, there are four agent types: households, firms, a government, and a central bank, each simulated by an LLM. Building on prior work (Gatti et al., 2011; Wolf et al., 2013; Dawid & Gatti, 2018; Li et al., 2024), our enriched framework expands the decision space considerably.
@@ 4.1 Households
Households are the fundamental units in our simulation and interacts with all other agent types and markets. Each household is initialized with a heterogeneous profile, including age, education, consumption preference, skill endowment. Details about these characteristics and initialization process can be found in Appendix D.4.
Each month, households receive two types of information: a personal report summarizing income, expenditures, and other status changes during the current period, and a citywide update containing goods prices, labor market conditions, housing availability, returns on investment and interest rates. Additionally, households are matched with job vacancies created by firms probabilistically.
@@ 4.2 Investment Pool and Firms
The pool chooses from a library of 44 synthetic firm templates, each of which produces a unique type of good, and selects a a geograhpic placement for the firm. The construction of these templates from real-world data is detailed in Appendix D.5. After a firm is established, all contributing households receive shares of the firm in proportion to their investment.
@@
Each firm in SimCity produces a single good. It hires households as workers, invests in capital to improve productivity, and transforms input goods to its specialized output, which is then sold in the market. Firms are instantiated from templates that specify skills requirements of job positions in the firm and the input-output relations governing its production process, where a firm may take possibly more than one input good for its production. The effectiveness of resource utilization (i.e. the ratio at which input goods are transformed into the output good) is determined by a firm-specific Cobb-Douglas production function (Cobb & Douglas, 1928):
@@
where $A$ is total factor productivity, $K$ the firm’s capital stock, $L$ the effective labor supply, defined as the amount of labor input adjusted for the match between job skill requirements and employees’ actual skills.
@@
More detailed description of firms can be found at Appendix C.3.
@@ 4.3 Government
where $\cdot]_{+}\equiv\max\{\cdot,0\}$. The total tax revenue collected by the government is the summation of taxes from all agents. The government uses this tax revenue in three ways: investing in the construction of public service buildings, which are modeled as government-owned firms, distributing it to households as a universal basic income (UBI), or reserving it.
@@ 4.4 Central Bank and Financial System
The central bank accepts deposits, provides loans, and implements monetary policy. In the first month of each year, it sets the policy interest rate according to a modified Taylor rule, which is a widely used monetary policy targeting rule (Galí & Gertler, 2007; Dawid & Gatti, 2018). Formally:
@@
where $\hat{r}$ is the policy rate set by the central bank, $r^{n}$ the long-run natural interest rate, $\pi^{t}$ the target inflation rate, $\pi$ the GDP inflation in the last period, $Y$ the actual output (GDP), and $Y^{n}$ the potential output as measured by a linear trend. Parameters $\alpha$ and $\beta$ capture the central bank’s responsiveness to inflation and output gap, respectively.
Intuitively, when inflation is high or output exceeds its long-term trend (signals of an overheated economy), the central bank raises interest rate to cool the economy and maintain monetary stability, and vice versa (Galí, 2015). To reflect the gradual adjustment observed in real-world monetary policy, we incorporate a smoothing term consistent with empirical evidence on persistent rate changes (Coibion & Gorodnichenko, 2012).
At each time step, deposits and loans accrue interest: deposit rate equals the policy interest rate, while loan rate equals the policy rate plus a fixed markup. Further details about the financial system are in Appendix C.5.
@@ 5 Experiments
Table 2: Macroeconomic agent-based modeling checklist.
Regularity AI-economist EconAgent LEN CATS Simcity(Ours)
Phillips Curve [^3] Short-run inverse relationship between the inflation rate and the unemployment rate. // $\checkmark$ $\times$ $\times$ $\checkmark$
Okun’s Law [^4] Inverse relationship between the change in the unemployment rate and the change in real GDP growth. // $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Beveridge Curve [^5] Job vacancy rate tends to fall when unemployment rises. // // $\times$ [^6] The LEN paper claims verified Beveridge Curve, but we are unable to reproduce with https://github.com/newwayland/baseline-economy. // $\checkmark$
Price Elasticity of Demand [^7] Quantity demanded is downward sloping in price, though the magnitude depends on the type of goods. // // // // $\checkmark$
Engel’s Law [^8] As household income rises, food’s spending share declines while the absolute amount increases. // // // // $\checkmark$
Investment Volatility [^9] Consumption is less volatile than GDP, while investment is more volatile than GDP over the business cycle. // // // // $\checkmark$
Price Stickiness [^10] Price has stickiness, i.e., firms don’t update price frequently. This is also often called “menu cost”. // // // // $\checkmark$
Legend: $\checkmark$: Verified; $\times$: Verified but inconsistent; //: Mechanism not available in the model.
We conduct experiments to evaluate SimCity’s ability to perform macroeconomic simulations and to display urban development dynamics. A key limitation of existing frameworks (Lengnick, 2013; Li et al., 2024) is the lack of a common basis for comparison. Prior studies often highlight a relatively small set of selected macroeconomic phenomena, making it difficult to gauge and compare the capacities of different simulation frameworks. To address this gap, we compile a checklist of canonical macroeconomic “stylized facts” from the broad economics literature (Blanchard & Fischer, 1989; Williamson, 2014; Òscar Jordà et al., 2017; Axtell & Farmer, 2025), many of which have previously been used to evaluate other ABM models. We then use this checklist to assess SimCity’s ability to reproduce established patterns and to simulate new ones under novel economic shocks. Taking advantage of SimCity’s rich framework, we aim to answer the following research questions:
@@
  • RQ2: Are emergent regularities robust across multiple simulations?
  • RQ3: How does SimCity grow during the move-in phase?
  • RQ4: To what extent can SimCity reflect against external shock?
@@ 5.0.1 Experiment Setup
We simulate an economy with a maximum of 200 households. Simulation proceeds in two phases. During the initial phase, which lasts 36 steps (3 years), new households continue to enter. During the second phase, which lasts 144 steps (12 years, 48 quarters), the population remains fixed and we observe the simulation under steady-state conditions. We use gpt-4o-mini for regular reasoning and gpt-4 as the vision language model for establishment decision-making. Both are provided by Azure OpenAI API [^11] https://portal.azure.com/ with default sampling parameter. We provide detailed examples of prompts and responses in Appendix E.3.
For comparison, we select two traditional ABMs with predetermined rules LEN (Lengnick, 2013) and CATS (Gatti et al., 2011), a deep multiagent reinforcement learning model AI-economist (Zheng et al., 2022), a LLM-based system EconAgent (Li et al., 2024) and real world data FRED from 1970 Q1 (Federal Reserve Economic Data) [^12] https://fred.stlouisfed.org/.
5.1 Emergence of Macroeconomic Phenomena (RQ1)
A central test of macroeconomic simulations is its ability to exhibit well-documented empirical regularities (“stylized facts”) observed in real economies. Leveraging the strong role-play capabilities of LLM agents, SimCity captures many of these regularities that traditional agent-based models have historically struggled to generate. Table 2 summarizes the key phenomena and compares SimCity’s performance with baseline models.

### v2

@@ 1 Introduction
The rapid development of Large Language Models (LLMs) has enabled multi-agent simulations of human societal activities across diverse scales and domains (Gao et al., 2024a). In these settings, autonomous LLM-powered agents interact with each other and with their environment. Existing work has examined not only general social simulacra (Park et al., 2023; Huang et al., 2025; Piao et al., 2025), but also domain-specific applications such as public administration crisis (Xiao et al., 2023), health policy (Hou et al., 2025), or deduction game (Xu et al., 2025). While these works demonstrate the social simulation capabilities of LLM-driven agents, we focus on their application to the simulation and evaluation of urban-style economic activities.
For the past two decades, the Dynamic Stochastic General Equilibrium (DSGE) framework has been the predominant paradigm for studying aggregate economic behavior (Blanchard, 2009; Glandon et al., 2023). While mathematically elegant, DSGE relies on the assumption of simplified representative agents solving explicit optimization problems, which restricts heterogeneity and richer behavioral dynamics (Sergi, 2018; Vines and Wills, 2020; Storm, 2021). Agent-based models (ABMs) offer a bottom-up alternative that accommodates heterogeneity but typically rely on hand-crafted decision rules (Dilaver et al., 2018). Advancing the ABM tradition with recent developments in LLMs, we propose SimCity, a multi-agent macroeconomic simulation framework that enables flexible, adaptive behavior with transparent natural language reasoning, and a virtual city environment that provides spatial context.
SimCity models the economy as the interaction of four types of agents: households, firms, a central bank, and a government. To incorporate realistic heterogeneity, households and firms are instantiated as families of agents with rich variation in preferences, abilities, and other background, whereas the central bank and government are modeled as single institutional agents, in line with the economic literature (Blanchard, 2025). Agents are implemented through an LLM-based module that integrates environmental observations, agent-specific traits, and structured memory to generate reasoning, planning, and decision-making. Following the classic agent–interaction–environment paradigm of multi-agent system design (Wooldridge, 2009), we construct a simulation environment that features a frictional labor market, heterogeneous goods markets, and core financial interactions. To provide spatial context for both agents and researchers, the environment is augmented with a visualized map (Figure 1).
@@
Experimental results show that SimCity successfully exhibits a range of classic macroeconomic phenomena, such as Okun’s law, the Phillips curve, the Beveridge curve, price elasticity of demand, and Engel’s curve (detailed in Section 5). It also simulates the dynamic expansion of a city, and facilitates the simulation of novel economic shocks that were difficult to study in previous frameworks.
@@
1. LLM-driven macroeconomic simulation with rich interactions. We integrate LLMs into economic simulation by modeling four distinct economic roles, households, firms, a central bank, and a government, as LLM-driven agents. This enables realistic agent heterogeneity and rich economic interactions.
@@
2 Related Work
DSGE and ABM represent two leading, and in many respects, contrasting approaches to macroeconomic modeling. DSGE features highly idealized decision-makers characterized by precise mathematical objective functions, rational expectations about future variables, and strict market-clearing equilibrium conditions (Stokey et al., 1989; Ljungqvist and Sargent, 2018). In contrast, ABM follows a bottom-up, simulation-driven methodology, where agents are endowed with rule-of-thumb behavior patterns and, in some cases, learning heuristics (Tesfatsion and Judd, 2006; Axtell and Farmer, 2025). Although DSGE remains the dominant approach in economics departments and policy institutions (Woodford, 2009; Negro et al., 2013; Chen et al., 2023), there has been a renewal of interest in ABMs, particularly following the Global Financial Crisis of 2008 and again after the COVID-19 crisis, as mainstream DSGE models struggled to account for these unprecedented shocks (Gatti et al., 2008; Stiglitz, 2018; Borsos et al., 2024). Our work is closer to the ABM tradition. However, we depart from conventional ABMs by replacing predetermined behavioral rules with LLM-powered agents capable of flexible reasoning and information processing in natural languages.
Li et al. (2024) introduces EconAgent as a recent attempt at LLM-driven macroeconomic simulation. Its framework restricts agent-environment interactions to two simplified variables: consumption share and work propensity, which constrains the range of economic activities that can be simulated. Our work extends this approach by incorporating other key economic elements, including firms as LLM-driven agents, heterogeneous goods markets, enriched financial activities, and a taxation system. Thus, our framework allows the exploration of a wider range of macroeconomic phenomena. Another contemporary work Mi et al. (2025), which build similar cluster of agents, is proposed as a testbed for performance evaluation while we focus on urban development and macroeconomic laws.
Simulator AI Econ. LEN CATS EconAgent SimCity
Households 10 100 100 200 1000
@@ 2 Related Work
VAT
Agents’ Types HG HB HFB HBG HFBG
Markets $\times$ L LGF $\times$ LGF
@@
Note: VAT: Value-Added Tax. H: Household, F: Firm, B: Bank, G: Gov. L: Labor, G: Goods, F: Financial.
Table 1: A comparison of agent-driven macroeconomic simulators.
Regularity AI- Econ. Econ Agent LEN CATS SimCity (Ours)
Phillips Curve / $\times$ $\times$
Okun’s Law /
Beveridge Curve / / $\times$ /
Price Elasticity of Demand / / / /
Engel’s Law / / / /
Investment Volatility / / / /
Price
Stickiness / / / /
Legend: ✓: Verified; $\times$: Inconsistent; /: Not available.
Note: The LEN paper claims verified the Beveridge Curve, but we are unable to reproduce it with https://github.com/newwayland/baseline-economy
Table 2: Checklist of macroeconomic regularities considered in this work.
@@ 3 Environment and Interaction
Our framework consists of three core layers: environment, interaction protocol, and agents. This section describes the environment and interaction of our system architecture.
@@ 3.1 Environment
On a high level, the simulation proceeds in two phases. In phase 1 (the move-in phase), new households with synthetic profiles, representing immigrants or newborns, are introduced into the SimCity environment until a predetermined maximum population is reached. Unlike the typical burn-in period used purely for model stabilization in ABMs (Lengnick, 2013), this phase is explicitly designed to study urban expansion, capturing how economic activity and spatial development evolve as the city grows.
In phase 2 (the development phase), the population is fixed. However, urban developments such as firm creation, investment, and spatial reallocation, continue to take place. We focus on this phase to study macroeconomic regularities, as the system operates in a quasi-steady state that is less susceptible to the transient noise present during the expansion period.
Each phase lasts some steps. Each simulation step represents one month. The order of events within each step is detailed in Appendix D.2. By design, agents can only access information from previous steps.
@@
We develop a web-based render module to visualize the urban expansion as shown in Figure 1. All buildings are displayed on the rendered map. The geographic placement of new firms is decided by a Visual-Language Model (VLM). For technical details of the module, please refer to Appendix D.3.
@@ 3.2 Interaction Protocol
We leverage the common-sense reasoning capabilities of LLMs to act as human-like, heterogeneous agents. Details about the prompts used are provided in Appendix E.1. Agents interact with the environment by means of function calling. The framework loads all the operations that the agent can execute and appends formatted function names along with their descriptions to the prompt. The LLMs will return the actions to be taken and their parameters in JSON format. The framework will execute after a verification. For detailed examples, see Appendix E.3.
@@ 4 Agents
This section presents an overview of the agents in SimCity. As illustrated in Figure 1, there are four agent types: households, firms, a government, and a central bank, each simulated by an LLM. Building on prior work (Gatti et al., 2011; Wolf et al., 2013; Dawid and Gatti, 2018; Li et al., 2024), our enriched framework expands the decision space considerably.
@@ 4.1 Households
Households are the fundamental units in our simulation and interact with all other agent types and markets. Each household is initialized with a heterogeneous profile, including age, education, consumption preference, and skill endowment. Details about these characteristics and the initialization process can be found in Appendix D.4.
Each month, households receive two types of information: a personal report summarizing income, expenditures, and other status changes during the current period, and a citywide update containing goods prices, labor market conditions, housing availability, returns on investment, and interest rates. Additionally, households may receive offers for vacancies based on their skill alignment.
@@ 4.2 Investment Pool and Firms
The pool chooses from a library of 44 synthetic firm templates, each of which produces a unique type of good, and selects a geographic placement for the firm. The construction of these templates from real-world data is detailed in Appendix D.5. After a firm is established, all contributing households receive shares of the firm in proportion to their investment.
@@
Each firm in SimCity produces a single type of good. It hires households as workers, invests in capital to improve productivity, and transforms input goods to its specialized output, which is then sold in the market. Firms are instantiated from templates that specify skill requirements of job positions in the firm and the input-output relations governing its production process, where a firm may take possibly more than one input good for its production. The effectiveness of resource utilization (i.e., the ratio at which input goods are transformed into the output good) is determined by a firm-specific Cobb-Douglas production function (Cobb and Douglas, 1928):
@@
where $A$ is total factor productivity, $K$ is the firm’s capital stock, $L$ is the effective labor supply, defined as the amount of labor input adjusted for the match between job skill requirements and employees’ actual skills.
@@
More detailed descriptions of firms can be found in Appendix C.3.
@@ 4.3 Government
where $\cdot]_{+}\equiv\max\{\cdot,0\}$. The total tax revenue collected by the government is the summation of taxes from all agents. The government uses this tax revenue in three ways: investing in the construction of public service buildings, which are modeled as government-owned firms, distributing it to households as a universal basic income (UBI), or reserving it for the future.
@@ 4.4 Central Bank and Financial System
The central bank accepts deposits, provides loans, and implements monetary policy. In the first month of each year, it sets the policy interest rate according to a modified Taylor rule, which is a widely used monetary policy targeting rule (Galí and Gertler, 2007; Dawid and Gatti, 2018). Formally:
@@
where $\hat{r}$ is the policy rate set by the central bank, $r^{n}$ is the long-run natural interest rate, $\pi^{t}$ is the target inflation rate, $\pi$ is the GDP inflation in the last period, $Y$ is the actual output (GDP), and $Y^{n}$ is the potential output as measured by a linear trend. Parameters $\alpha$ and $\beta$ capture the central bank’s responsiveness to inflation and output gap, respectively.
Intuitively, when inflation is high, or output exceeds its long-term trend (signals of an overheated economy), the central bank raises interest rates to cool the economy and maintain monetary stability, and vice versa (Galí, 2015). To reflect the gradual adjustment observed in real-world monetary policy, we incorporate a smoothing term consistent with empirical evidence on persistent rate changes (Coibion and Gorodnichenko, 2012).
At each time step, deposits and loans accrue interest: the deposit rate equals the policy interest rate, while the loan rate equals the policy rate plus a fixed markup. Further details about the financial system are in Appendix C.5.
@@ 5 Experiments
We conduct experiments to evaluate SimCity’s ability to perform macroeconomic simulations and to display urban development dynamics. A key limitation of existing frameworks (Lengnick, 2013; Li et al., 2024) is the lack of a common basis for comparison. Prior studies often highlight a relatively small set of selected macroeconomic phenomena, making it difficult to gauge and compare the capacities of different simulation frameworks.
To address this gap, we compile a checklist of canonical macroeconomic “stylized facts” from the broad economics literature (Blanchard and Fischer, 1989; Williamson, 2014; Jordà et al., 2017; Axtell and Farmer, 2025), many of which have previously been used to evaluate other ABM models. We then use this checklist to assess SimCity’s ability to reproduce established patterns and to simulate new ones under novel economic shocks. Taking advantage of SimCity’s rich framework, we aim to answer the following research questions:
@@
  • RQ2: Are emergent regularities robust across multiple independent simulations?
  • RQ3: Can SimCity exhibit realistic spatial patterns during the move-in and urban expansion phase?
  • RQ4: To what extent can SimCity respond to external shocks?
@@ 5.0.1 Experiment Setup
We simulate an economy with a maximum of 1000 households. Simulation proceeds in two phases. During the initial phase, which lasts 36 steps (3 years), new households gradually populate the city. During the second phase, which lasts 144 steps (12 years, 48 quarters), the population remains fixed and we observe the simulation under steady-state conditions. We use gpt-4o-mini for regular reasoning and gpt-4 as the vision language model for establishment decision-making. Both are provided by Azure OpenAI API [^1] https://portal.azure.com/ with default sampling parameter. We provide detailed examples of prompts and responses in Appendix E.3.
For comparison, we select two traditional ABMs with predetermined rules LEN (Lengnick, 2013) and CATS (Gatti et al., 2011), a deep multiagent reinforcement learning model AI-economist (Zheng et al., 2022), a LLM-based system EconAgent (Li et al., 2024), and real-world data FRED from 1970 Q1 (Federal Reserve Economic Data) [^2] https://fred.stlouisfed.org/.
Figure 2: Emergence of the Phillips Curve and Okun’s Law in SimCity simulations. $r$-value is the Pearson correlation coefficient, and $p$-value indicates the statistical significance of it.
Figure 3: Beveridge Curve and other macroeconomic emergence from SimCity.
5.1 Macroeconomic Emergence (RQ1)
A central test of macroeconomic simulations is its ability to exhibit well-documented empirical regularities (“stylized facts”) observed in real-world economies. Leveraging the strong role-play capabilities of LLM agents, SimCity captures many of these regularities that traditional agent-based models have historically struggled to generate. Table 2 summarizes the key phenomena and compares SimCity’s performance with baseline models.